# A multi-process imaging of nuclear modifications on parton distributions in proton-nucleus collisions

### v1

@@ III A multi-process imaging of $r^{\textrm{A}}_{i}$
Figure 6: Nuclear modification factors $R_{pA}$ as functions of $X_{B}$, calculated at NLO for reorganized cross section of $Z+\!c\!-\!$jet production (blue triangles) and for combined observable $C_{\textrm{charm}}(X_{B})$ (red discs). Predictions using EPPS21, nCTEQ15, and TUJU19 nuclear PDFs are displayed in left, middle, and right panels, respectively. Underlying $r^{\textrm{A}}_{i}(x,Q^{2})$ factors for light quarks, gluons and charm quarks from corresponding nuclear PDFs at $Q\!=\!M_{Z}/2$ are included in each panel for comparison. Bottom panels illustrate differences between each result and $r^{\textrm{A}}_{i}$ for charm quark. NLO calculations are conducted at scale $\mu_{0}\!=\!M_{Z}/2$ for $Z$-boson and at $\mu_{0}\!=\!(M_{V\!J}\!+\!p_{T,avg})/4$ for $Z+\!(c)$jet. Uncertainties in the results, evaluated by varying scale between $\mu_{0}/2$ and $2\mu_{0}$, are represented by shaded bands.
@@
Following a similar methodology, we now investigate the production of $Z\!+\!c$-jet in $p$A collisions, a process that is sensitive to the nuclear modifications of both gluon and charm quark distributions.
@@
The $R_{pA}$ for the reorganized $Z\!+\!c$-jet cross section, as defined in Eq. (4), is calculated at NLO using three sets of nuclear PDFs. These results are depicted as blue triangles in Fig. 6, alongside the corresponding $r^{\textrm{A}}_{i}(x_{\textrm{Pb}})$ curves for light-quark (solid), gluon (dotted-dashed), and charm (dotted) distributions at $Q=M_{Z}/2$, included for comparative analysis. To maintain consistency in probing scales with the aforementioned processes, we employ kinematic cuts of $15066<M_{V\!J}<$ 150 GeV and $6020<p_{T,avg}<$ 60 GeV in the calculations of the reorganized $Z\!+\!c$-jet cross section. As illustrated in Fig. 6, the $R_{pA}$ triangles closely align with the $r^{\textrm{A}}_{i}$ curves for both gluon and charm quark distributions, which are predicted to be similar across the three parametrizations. Additionally, the differences between each result and the corresponding $r^{\textrm{A}}_{i}$ for charm quark are presented in the bottom panels of Fig. 6 for further comparison.
@@
Using the definition of $C_{\textrm{charm}}(X_{B})$, we calculate its nuclear modification factor $R_{pA}(X_{B})$ at NLO for three nuclear PDFs and present the results as red discs in Fig. 6. As anticipated, since the gluon and charm distributions exhibit similar modifications in the three nuclear PDFs, the $R_{pA}$ factors for the combined $C_{\textrm{charm}}$ show only minor deviations from those for the individual $Z\!+\!c$-jet cross section (blue triangles). Moreover, they slightly better reproduce the $r^{\textrm{A}}_{i}(x_{\textrm{Pb}})$ curves for the charm quark distribution. This demonstrates that the $R_{pA}(X_{B})$ of $C_{\textrm{charm}}$ can effectively serve as an image of the $r^{\textrm{A}}_{i}$ for the charm quark distribution.
@@ IV SUMMARY AND DISCUSSION
Nuclear modifications to parton distribution functions are primarily inferred through global analyses of a wide range of data within the framework of collinear factorization. However, determining the explicit dependence of the modification factors $r^{\textrm{A}}_{i}(x,Q^{2})$ on the variables $x$, $Q^{2}$, and parton species $i$ remains challenging. This is partly due to their complex relationships with observables and the inherent uncertainties in the data. Although PDFs and their modification factors $r^{\textrm{A}}_{i}(x,Q^{2})$ are not directly measurable, developing theoretically-assisted observables can not only enhance the efficiency of global extraction but also provide deeper insights into the underlying physics.
@@
In this work, we introduce a series of novel observables in proton-nucleus collisions at the LHC, designed to effectively scan the underlying nuclear modification factors $r^{\textrm{A}}_{i}(x,Q^{2})$. Specifically, by combining the reorganized cross sections of $Z$-boson production, $Z$+jet production, and $Z+c$-jet production, we separately amplify signals from light-quark, gluon, and heavy-flavor (charm) distributions in nuclei. This approach allows us to effectively image the $r^{\textrm{A}}_{i}(x,Q^{2})$ for specific parton species, serving as an analogy to the measurement of nuclear modifications on structure functions in deep-inelastic scattering (DIS).
@@
The feasibility of this method is validated through perturbative calculations at next-to-leading order using three parametrizations of nuclear PDFs. Measurements of these observables are expected to yield valuable insights into the partonic structures within nuclei. While the current precision of LHC measurements may not yet suffice to achieve a perfect image of $r^{\textrm{A}}_{i}(x,Q^{2})$, the method proposed here opens new possibilities for optimizing future measurements. Notably, this imaging technique can be generalized to study other processes at the LHC, such as the Drell-Yan process Reimer:2007zza ; CMS:2021ynu , dijet production CMS:2018jpl , and $\gamma+$jet production Neufeld:2010fj ; Ma:2018tjv . Its future applications in the EIC program, e.g., for dijet production Guzey:2020zza ; Zhang:2021tcc , are also of significant interest. Importantly, this approach enables systematic comparisons of nuclear modifications across various processes, spanning from $p$A to $e$A collisions.
@@ Acknowledgements.
The authors would like to thank G. Moore and X.-N. Wang for helpful discussions. This research was supported in part by Guangdong Major Project of Basic and Applied Basic Research No. 2020B0301030008, by Guangdong Basic and Applied Basic Research Foundation under Project No. 2022A1515110392, and by Natural Science Foundation of China (NSFC) under Project Nos. 11935007 and 11805167.

### v2

@@ III A multi-process imaging of $r^{\textrm{A}}_{i}$
Figure 6: Nuclear modification factors $R_{pA}$ as functions of $X_{B}$, calculated at NLO for reorganized cross section of Z+cZ+\!c\!-\!jet production (blue triangles) and for combined observable $C_{\textrm{charm}}(X_{B})$ (red discs). Predictions using EPPS21, nCTEQ15, and TUJU19 nuclear PDFs are displayed in left, middle, and right panels, respectively. Underlying $r^{\textrm{A}}_{i}(x,Q^{2})$ factors for light quarks, gluons and charm quarks from corresponding nuclear PDFs at $Q\!=\!M_{Z}/2$ are included in each panel for comparison. Bottom panels illustrate differences between each result and $r^{\textrm{A}}_{i}$ for charm quark. NLO calculations are conducted at scale $\mu_{0}\!=\!M_{Z}/2$ for $Z$-boson and at $\mu_{0}\!=\!(M_{V\!J}\!+\!p_{T,avg})/4$ for Z+(c)Z+\!(c)jet. Uncertainties in the results, evaluated by varying scale between $\mu_{0}/2$ and $2\mu_{0}$, are represented by shaded bands.
In the calculations, the renormalization and factorization scales are set to $\mu_{0}=(M_{V\!J}+p_{T,avg})/4$ for Z+Z+jet production, while for $Z$-boson production, they are set to $\mu_{0}=M_{Z}/2$, consistent with the previous section. For Z+Z+jet production, the variations in $R_{pA}$ with $\mu_{0}$ scaled by a factor of 2, represented as shaded bands around the squares in Fig. 5, are as small as those in $Z$-boson production. To illustrate the scale dependence for the $R_{pA}$ of the combined observable $C_{g}$, we assume that the choices of $\mu_{0}$ in the two processes are independent. Given that the ratio between $\mu^{2}_{0}$ in the two processes can vary by a factor of 16, the scale dependence of the combined observable generally appears stronger than that of a single observable, as shown in Fig. 5. The $R_{pA}$ of $C_{g}$ effectively images the $r^{\textrm{A}}_{i}$ for the gluon distribution up to a moderate value of $x_{\textrm{Pb}}$ ($\sim$0.1). For larger values of $x_{\textrm{Pb}}$, the uncertainties become significant, related to the rapidly diminishing contribution from gluons.
It should be mentioned that, as an input in the definition of $C_{g}(X_{B})$, the coefficient $\kappa_{1}(X_{B})$ itself depends on the theoretical setting in its determination. A possible impact of the free-proton PDF (e.g. scheme and parametrization) mainly arises from the relative compositions of various flavors in proton. To test the effects from proton PDF parametrization, we calculate the contribution fraction for certain nuclear parton flavors (quark/gluon) with three parametrizations of free-proton PDFs (CT18 Hou:2019efy , MMHT Harland-Lang:2019pla and NNPDF Cruz-Martinez:2024cbz , see the Appendix A), and observe negligible distinctions among these results in most of the studied kinematic regions in this work. Besides, the PDF/factorization scheme may also have an impact. For example, at NLO and beyond the division between quark- and gluon-initiated channels is factorization-scheme dependent. Generally, the differences between the predictions with different schemes may become reduced when higher-order contributions are included. A quantitative study of the impact of the PDF scheme/factorization is beyond the scope of this work. Note that it is impossible to directly measure $r^{\textrm{A}}_{i}(x,Q^{2})$ or completely purify the signal from a certain flavor. Therefore, the observable $C_{g}(X_{B})$ is designed to establish an approximate and effective mapping to the underlying $r^{\textrm{A}}_{i}(x,Q^{2})$ factors for gluon. In this work, the effectiveness of such observables is examined by NLO calculations with scale variations, and a nice reflection to the object $r^{\textrm{A}}_{i}(x,Q^{2})$ is observed under the current scheme  Campbell:1999ah ; Campbell:2011bn ; Campbell:2015qma ; Campbell:2021vlt ; Hou:2019efy .
@@
Following a similar methodology, we now investigate the production of Z+cZ\!+\!c-jet in ppA collisions, a process that is sensitive to the nuclear modifications of both gluon and charm quark distributions.
@@
The $R_{pA}$ for the reorganized Z+cZ\!+\!c-jet cross section, as defined in Eq. (4), is calculated at NLO using three sets of nuclear PDFs. These results are depicted as blue triangles in Fig. 6, alongside the corresponding $r^{\textrm{A}}_{i}(x_{\textrm{Pb}})$ curves for light-quark (solid), gluon (dotted-dashed), and charm (dotted) distributions at $Q=M_{Z}/2$, included for comparative analysis. To maintain consistency in probing scales with the aforementioned processes, we employ kinematic cuts of $66<M_{V\!J}<150$ GeV and $20<p_{T,avg}<60$ GeV in the calculations of the reorganized Z+cZ\!+\!c-jet cross section. As illustrated in Fig. 6, the $R_{pA}$ triangles closely align with the $r^{\textrm{A}}_{i}$ curves for both gluon and charm quark distributions, which are predicted to be similar across the three parametrizations. Additionally, the differences between each result and the corresponding $r^{\textrm{A}}_{i}$ for charm quark are presented in the bottom panels of Fig. 6 for further comparison.
@@
Using the definition of $C_{\textrm{charm}}(X_{B})$, we calculate its nuclear modification factor $R_{pA}(X_{B})$ at NLO for three nuclear PDFs and present the results as red discs in Fig. 6. As anticipated, since the gluon and charm distributions exhibit similar modifications in the three nuclear PDFs, the $R_{pA}$ factors for the combined $C_{\textrm{charm}}$ show only minor deviations from those for the individual Z+cZ\!+\!c-jet cross section (blue triangles). Moreover, they slightly better reproduce the $r^{\textrm{A}}_{i}(x_{\textrm{Pb}})$ curves for the charm quark distribution. This demonstrates that the $R_{pA}(X_{B})$ of $C_{\textrm{charm}}$ can effectively serve as an image of the $r^{\textrm{A}}_{i}$ for the charm quark distribution.
@@ IV SUMMARY AND DISCUSSION
Nuclear modifications to parton distribution functions are primarily inferred through global analyses of a wide range of data within the framework of collinear factorization. However, determining the explicit dependence of the modification factors $r^{\textrm{A}}_{i}(x,Q^{2})$ on the variables $x$, $Q^{2}$, and parton species $i$ remains challenging. This is partly due to their complex relationships with observables, the inherent uncertainties in both the data and predictions, and the parametrization dependence in global extractions Klasen:2023uqj . Besides continuous improvements in both experimental and theoretical precision, introducing observables with better disentangled contributions from PDF variables may help mitigate this situation.
@@
In this work, we introduce a series of novel observables in proton-nucleus collisions at the LHC, designed to establish an approximate and effective mapping to the underlying nuclear modification factors $r^{\textrm{A}}_{i}(x,Q^{2})$. Specifically, by combining the reorganized cross sections of $Z$-boson production, ZZ+jet production, and $Z+c$-jet production, we separately purify signals from light-quark, gluon, and heavy-flavor (charm) distributions in nuclei. This approach allows us to effectively image the $r^{\textrm{A}}_{i}(x,Q^{2})$ for specific parton species, serving as an analogy to the measurement of nuclear modifications on structure functions in deep-inelastic scattering (DIS). Such imaging observables are expected to significantly enhance the impact of LHC data by providing more direct and effective constraints on designing the parametrization forms of flavor-separated nuclear modifications, which can facilitate the global nPDF fit. These intuitive images of $r^{\textrm{A}}_{i}(x,Q^{2})$ can also help study the non-perturbative dynamics and binding mechanisms in nuclei Alekhin:2022tip ; Alekhin:2022uwc ; Cloet:2012td ; CLAS:2019vsb ; Cocuzza:2021rfn ; Dalal:2022zkg ; nCTEQ:2023cpo .
@@
The feasibility of this method is validated through perturbative calculations at next-to-leading order using three parametrizations of nuclear PDFs. While the current precision of LHC measurements may not yet suffice to achieve a perfect image of $r^{\textrm{A}}_{i}(x,Q^{2})$, the method proposed here opens new possibilities for optimizing future measurements.
It should be emphasized that, both PDFs and their modification factors $r^{\textrm{A}}_{i}(x,Q^{2})$ are not directly measurable, although this work demonstrates the ability of the imaging observables to reflect the underlying $r^{\textrm{A}}_{i}(x,Q^{2})$ factors. In the context of global PDF extraction Dulat:2015mca ; Segarra:2020gtj ; Hirai:2007sx ; Eskola:2009uj ; deFlorian:2011fp ; Kovarik:2015cma ; Wang:2016mzo ; Walt:2019slu ; Khanpour:2020zyu ; Eskola:2021nhw ; AbdulKhalek:2022fyi , utilizing the data with better disentangled contributions from various values of $i$ and $x$ may help reduce the correlations and degeneracy of the fitting parameters and streamline uncertainty propagation analysis, resulting in enhanced computational efficiency and stability in the global extraction. Compared to traditional observables, the proposed imaging observables can provide somewhat preprocessed data with reduced entanglement of the contributions from PDF variables. Nevertheless, if such imaging observables share the common data source with the traditional measurements used in the analysis, we do not expect any statistically significant improvement in fitting the imaging observables relative to the analysis with usual measurements.
It is also noteworthy that, this imaging technique can be generalized to study other processes at the LHC, such as the Drell-Yan process Reimer:2007zza ; CMS:2021ynu , dijet production CMS:2018jpl , and $\gamma+$jet production Neufeld:2010fj ; Ma:2018tjv . Its future applications in the EIC program, e.g., for dijet production Guzey:2020zza ; Zhang:2021tcc , are also of significant interest. Importantly, this approach can help establish a framework for analysis and interpretation of the commonalities and distinctions in nuclear modification effects across diverse processes, spanning from ppA to eeA collisions. Furthermore, the study of nuclear effects at higher-twist, e.g. the parton multiple scattering Kang:2013ufa ; Ru:2019qvz ; LHCb:2021vww , may also benefit from such systematic comparisons, in which the leading-twist effects across various processes are effectively unified or mapped.
Figure 8: Contribution fraction of certain nuclear parton flavors in Z+Z+jet (left) and $Z+c$-jet (right) productions with three parametrizations of free-proton PDFs: CT18 Hou:2019efy , MMHT Harland-Lang:2019pla and NNPDF4.0 Cruz-Martinez:2024cbz .
@@ Acknowledgements.
The authors would like to thank G. Moore and X.-N. Wang for helpful discussions. This research was supported in part by National Natural Science Foundation of China (NSFC) under Project No. 12535010 and by Guangdong Basic and Applied Basic Research Foundation under Project No.2022A1515110392.
Appendix A A test of impact from proton PDF parametrization
To test the impact of the uncertainties in the relative compositions of various flavors in proton from PDF parametrization, we calculate the contribution fraction for certain nuclear parton flavors in Z+Z+jet and $Z+c$-jet productions with three parametrizations of free-proton PDFs (CT18 Hou:2019efy , MMHT Harland-Lang:2019pla and NNPDF Cruz-Martinez:2024cbz ). Results are plotted in Fig.8, and small distinctions among these results can be observed.